# A Computational Theory for Efficient Model Evaluation with Causal Guarantees

### v3

@@ Appendix B Evaluation scenes
The samples number of scenes Alert, Withdrawal, Higgs, Climate, Grid, Insurance, TERECO, Nuscale, kin8nm, Calculus, Ad, and A-share trade are listed in table 4. All the computation is on the one 13-inch MacBook Pro 2020 with Apple M1 chip and 16GB memory in 2 days.
@@
B.1 Scenes to test evaluation model
For 11 scenes to test evaluation model, 2000 subjects was randomly sampling from heterogeneous subject space. 20% data was used to build the real systems to generate true evaluation metrics and other 80% data was used to get the proxy metrics of a subject. When the evaluation samples are collected from the built real systems, 20% samples are used for evaluation model testing, and other 80% samples are for evaluation model learning. The experiment is performed 30 times. All the null value was removed from the raw data in preprocessing.
@@ B.1 Scenes to test evaluation model
B.2 Scenes to test conditional evaluation model
We collected the A-share data of China from 2024-8 to 2024-11 (89 days) by an open source tool AkShare. The collected day-wise features of stocks are “Profit Ratio”, “Average Cost”, “90 Cost Low”, “90 Cost High”, “90 Episode Medium”, “70 Cost Low”, “70 Cost High”, “70 Episode Medium”, “Closing”, “Opening”, “Highest”, “Lowest”, “Transaction Volume”, “Transaction Amount”, “Amplitude”. The strategy is to buy 1 hand, sell 1 hand, or hold by the subject following the open price of the market. The slippage and commission is not considered in this work. If the stock was limit up (limit down), then we can only hold even the model decision is to buy (sell). The subject’s input is the feature of the stock last day and output is the decision this day. The subject will be deployed and run for 1 time slot (10 trading days, about 2 weeks), and the evaluation condition is feature of the stock in the last time slot. At the end of deployment in a time slot, we will sell all the hold stocks. When the end of the slot is limit down at open, we assume that it will be sold by 90 percents of open price in future.
In order to simulate the direct effect from the subject itself to the sensitive market, we add a float model which input is the vectorized subjects and output is float rate of open price ranged from -0.1% to +0.1%.
We use data of 200 stocks with minimal ID from all 3214 stocks due to our computational limit. The start of train day is from day 10 to day 41, the start of test day is from 41 to 61. The sampling number of subject is 200. The experiment was performed 5 times to reduce the influence of randomness. The proxy metric is the last 10 days’ RoI.
Appendix C Assumption assessment
Table LABEL:tab:assessregression and table LABEL:tab:assessclass are assumption assessment results of 7 kinds of decision models for continuous and discrete target respectively. From the tables, we can see that the IIDE assumption (IID check and ID check) and unbiased evaluation model assumption (Bias Check and GroupBias Check) was not violated obviously with a high probability (>0.8) except the cases that using MLP and SVR as base regression learners of HetEM.
Table 5: The ratio that the p-value of test is larger than or equal to 0.05 in the 30 repeated experiments for regression base learners. The name of each row is “BaseLearnerName”-“TestName”-“MetricName”. The name of each column is the scene name.
@@ Appendix C Assumption assessment
Linear-IID-RMSE 0.93 0.87 0.97 0.9 0.97
Linear-IID-$R^{2}$ 0.93 1.0 0.9 0.93 0.83
Linear-ID-RMSE 1.0 0.47 0.87 1.0 0.93
Linear-ID-$R^{2}$ 0.97 0.43 0.93 0.97 0.87
Linear-Bias-RMSE 0.9 0.43 0.77 0.8 0.77
Linear-Bias-$R^{2}$ 0.9 0.33 0.8 0.77 0.77
Linear-GroupBias-RMSE 0.92 0.85 0.87 0.87 0.85
Linear-GroupBias-$R^{2}$ 0.90 0.85 0.86 0.84 0.89
MLP-IID-RMSE 1.0 0.93 0.83 1.0 0.9
MLP-IID-$R^{2}$ 0.9 0.97 0.93 0.97 0.97
MLP-ID-RMSE 0.73 0.37 0.87 0.73 0.77
MLP-ID-$R^{2}$ 0.9 0.3 0.77 0.77 0.67
MLP-Bias-RMSE 0.5 0.26 0.57 0.57 0.5
MLP-Bias-$R^{2}$ 0.57 0.17 0.53 0.57 0.57
MLP-GroupBias-RMSE 0.68 0.40 0.86 0.70 0.72
MLP-GroupBias-$R^{2}$ 0.70 0.31 0.71 0.67 0.68
SVR-IID-RMSE 0.97 0.97 0.97 1.0 0.87
SVR-IID-$R^{2}$ 0.93 0.9 1.0 0.9 0.97
SVR-ID-RMSE 0.0 0.0 0.27 0.4 0.1
SVR-ID-$R^{2}$ 0.0 0.0 0.0 0.2 0.0
SVR-Bias-RMSE 0.0 0.0 0.1 0.17 0.03
SVR-Bias-$R^{2}$ 0.0 0.23 0.0 0.03 0.0
SVR-GroupBias-RMSE 0.0 0.01 0.19 0.34 0.11
SVR-GroupBias-$R^{2}$ 0.0 0.44 0.00 0.19 0.02
RF-IID-RMSE 0.97 0.93 0.97 0.97 0.93
RF-IID-$R^{2}$ 1.0 0.97 0.97 1.0 0.97
RF-ID-RMSE 0.93 0.87 0.9 0.97 0.83
RF-ID-$R^{2}$ 0.93 0.93 0.93 0.9 0.73
RF-Bias-RMSE 0.93 0.57 0.9 0.9 0.8
RF-Bias-$R^{2}$ 0.83 0.63 0.93 0.87 0.8
RF-GroupBias-RMSE 0.87 0.85 0.91 0.90 0.86
RF-GroupBias-$R^{2}$ 0.89 0.86 0.91 0.88 0.86
LGBM-IID-RMSE 0.87 0.93 0.97 1.0 0.93
LGBM-IID-$R^{2}$ 1.0 0.97 0.97 0.93 0.97
LGBM-ID-RMSE 0.93 0.9 0.9 0.97 0.83
LGBM-ID-$R^{2}$ 0.9 0.93 0.83 0.97 0.93
LGBM-Bias-RMSE 0.9 0.7 0.8 0.9 0.77
LGBM-Bias-$R^{2}$ 0.93 0.7 0.77 0.9 0.83
LGBM-GroupBias-RMSE 0.89 0.85 0.90 0.88 0.85
LGBM-GroupBias-$R^{2}$ 0.89 0.86 0.97 0.88 0.87
XGBoost-IID-RMSE 1.0 0.93 0.93 1.0 0.87
XGBoost-IID-$R^{2}$ 0.93 0.93 0.93 0.97 0.93
XGBoost-ID-RMSE 0.93 0.9 1.0 0.93 0.97
XGBoost-ID-$R^{2}$ 0.93 0.9 0.93 0.93 0.87
XGBoost-Bias-RMSE 0.97 0.53 0.77 0.87 0.87
XGBoost-Bias-$R^{2}$ 0.83 0.5 0.9 0.87 0.83
XGBoost-GroupBias-RMSE 0.92 0.84 0.83 0.86 0.88
XGBoost-GroupBias-$R^{2}$ 0.87 0.82 0.90 0.84 0.88
CatBoost-IID-RMSE 0.87 0.97 0.97 1.0 1.0
CatBoost-IID-$R^{2}$ 0.87 0.97 0.93 0.93 1.0
CatBoost-ID-RMSE 0.93 0.87 0.93 0.97 0.97
CatBoost-ID-$R^{2}$ 0.97 0.9 0.93 0.97 1.0
CatBoost-Bias-RMSE 0.93 0.6 0.97 0.8 0.83
CatBoost-Bias-$R^{2}$ 0.93 0.63 0.93 0.8 0.87
CatBoost-GroupBias-RMSE 0.90 0.90 0.92 0.87 0.87
CatBoost-GroupBias-$R^{2}$ 0.90 0.87 0.90 0.86 0.88
Table 6: The ratio that the p-value of test is larger than or equal to 0.05 in the 30 repeated experiments for classification base learners. The name of each row is “BaseLearnerName”-“TestName”-“MetricName”. The name of each column is the scene name.
@@
Linear-IID-ROCAUC 0.96 0.93 0.93 0.93 1.0 0.97
Linear-IID-ACC 0.93 0.97 0.9 0.9 0.97 0.97
Linear-ID-ROCAUC 1.0 0.97 0.87 0.9 1.0 0.97
Linear-ID-ACC 0.93 0.97 0.87 0.93 0.93 1.0
Linear-Bias-ROCAUC 0.77 0.83 0.83 0.9 0.9 0.9
Linear-Bias-ACC 0.83 0.87 0.87 0.8 0.83 0.93
Linear-GroupBias-ROCAUC 0.83 0.88 0.87 0.88 0.91 0.87
Linear-GroupBias-ACC 0.87 0.89 0.87 0.88 0.87 0.89
MLP-IID-ROCAUC 0.97 1.0 0.97 0.97 0.93 0.93
MLP-IID-ACC 0.93 0.93 0.97 0.9 0.97 0.97
MLP-ID-ROCAUC 0.8 0.87 0.77 0.87 0.87 0.7
MLP-ID-ACC 0.8 0.83 0.83 0.8 0.73 0.8
MLP-Bias-ROCAUC 0.43 0.57 0.53 0.53 0.63 0.47
MLP-Bias-ACC 0.67 0.6 0.57 0.57 0.43 0.6
MLP-GroupBias-ROCAUC 0.70 0.76 0.71 0.73 0.77 0.69
MLP-GroupBias-ACC 0.77 0.73 0.77 0.74 0.63 0.73
SVM-IID-ROCAUC 0.97 0.97 0.93 0.97 0.97 1.0
SVM-IID-ACC 0.97 0.97 0.97 0.97 1.0 1.0
SVM-ID-ROCAUC 0.73 0.5 0.9 0.6 0.83 0.8
SVM-ID-ACC 0.9 0.67 0.37 0.97 0.97 0.97
SVM-Bias-ROCAUC 0.4 0.27 0.1 0.37 0.53 0.47
SVM-Bias-ACC 0.97 0.37 0.43 0.93 0.8 0.93
SVM-GroupBias-ROCAUC 0.63 0.48 0.36 0.52 0.69 0.66
SVM-GroupBias-ACC 0.90 0.56 0.62 0.90 0.85 0.86
RF-IID-ROCAUC 0.97 0.97 0.9 0.97 0.97 0.97
RF-IID-ACC 0.93 0.97 0.87 0.93 0.9 0.9
RF-ID-ROCAUC 0.97 0.87 0.9 0.9 0.83 0.97
RF-ID-ACC 1.0 0.87 0.9 0.97 1.0 0.87
RF-Bias-ROCAUC 0.73 0.9 0.77 0.87 0.93 0.87
RF-Bias-ACC 0.9 0.93 0.8 0.97 0.8 0.7
RF-GroupBias-ROCAUC 0.85 0.87 0.84 0.9 0.91 0.86
RF-GroupBias-ACC 0.87 0.89 0.85 0.89 0.88 0.85
LGBM-IID-ROCAUC 1.0 0.97 0.97 0.97 0.87 1.0
LGBM-IID-ACC 0.93 0.97 0.93 0.9 1.0 0.9
LGBM-ID-ROCAUC 1.0 0.87 0.93 0.93 0.87 0.93
LGBM-ID-ACC 0.97 0.83 0.97 1.0 0.97 0.97
LGBM-Bias-ROCAUC 0.7 0.93 0.83 0.73 0.97 0.9
LGBM-Bias-ACC 0.83 0.9 0.77 0.83 0.8 0.8
LGBM-GroupBias-ROCAUC 0.81 0.88 0.85 0.89 0.91 0.86
LGBM-GroupBias-ACC 0.88 0.88 0.84 0.90 0.86 0.85
XGBoost-IID-ROCAUC 1.0 0.97 1.0 0.8 0.93 1.0
XGBoost-IID-ACC 0.93 1.0 0.97 1.0 0.93 0.87
XGBoost-ID-ROCAUC 0.9 0.83 0.9 0.9 0.87 1.0
XGBoost-ID-ACC 0.97 0.87 0.93 0.97 0.93 0.97
XGBoost-Bias-ROCAUC 0.83 0.97 0.83 0.9 0.83 0.83
XGBoost-Bias-ACC 0.87 0.93 0.9 0.83 0.8 0.87
XGBoost-GroupBias-ROCAUC 0.88 0.89 0.90 0.90 0.87 0.89
XGBoost-GroupBias-ACC 0.88 0.89 0.88 0.88 0.89 0.88
CatBoost-IID-ROCAUC 0.93 0.8 0.9 0.97 0.93 0.9
CatBoost-IID-ACC 1.0 0.9 1.0 0.87 1.0 0.97
CatBoost-ID-ROCAUC 0.97 0.93 0.93 0.93 0.87 0.97
CatBoost-ID-ACC 0.97 0.93 0.83 0.97 0.97 0.97
CatBoost-Bias-ROCAUC 0.83 0.87 0.8 0.8 0.9 0.87
CatBoost-Bias-ACC 0.87 0.9 0.93 0.87 0.8 0.8
CatBoost-GroupBias-ROCAUC 0.86 0.89 0.88 0.86 0.89 0.87
CatBoost-GroupBias-ACC 0.85 0.89 0.89 0.90 0.86 0.86
@@
We also visualization the QQ plot of the baseline evaluation model in figure 5 and the assumption assessment result of our evaluation model in table 7. From the visualization, the error distribution of baseline is close to normal distribution. Although the HetEM-Linear and HetEM-CatBoost does not pass our ID check and Bias check, the dramatically error reduction of our evaluation model is still believed as valuable while it can be not used to calculate the upper bound of the 2 evaluation model directly in this scene.
@@
Linear-IID 1.0
Linear-ID 0.0
Linear-Bias 0.0
CatBoost-IID 1.0
CatBoost-ID 0.0
CatBoost-Bias 0.0
Table 7: The ratio that p-value of test is larger than or equal to 0.05 in the 5 repeated experiments for A-share trade scene. The name of each row is “BaseLearnerName”-“TestName”-“MetricName”. The name of each column is the scene name.
Figure 5: QQ plot of evaluation error of Baseline (Last10DayRoI).
Appendix D Hyper-parameters of evaluation model

### v4

@@ Appendix D Evaluation scenes
The samples number of scenes Alert, Withdrawal, Higgs, Climate, Grid, Insurance, TERECO, Nuscale, kin8nm, Calculus, Ad, and A-share trade are listed in table 5. The 7 kinds of base learners we used is Linear, MLP, SVM/SVR, RF (Breiman [2001]), LGBM (Aziz et al. [2022]), XGBoost (Chen and Guestrin [2016]), and CatBoost (Prokhorenkova et al. [2018]). All the computation is on the one 13-inch MacBook Pro 2020 with Apple M1 chip and 16GB memory in 2 days.
@@
D.1 Scenes to test evaluation model
For 11 scenes to test evaluation model, 2000 agents was randomly sampling from heterogeneous agent space. 20% data was used to build the real systems to generate true evaluation metrics and other 80% data was used to get the proxy metrics of a agent. When the evaluation samples are collected from the built real systems, 20% samples are used for evaluation model testing, and other 80% samples are for evaluation model learning. The experiment is performed 30 times. All the null value was removed from the raw data in preprocessing. The classical features are encoded by the label encoder to reduce the computation. The text classical features are encoded by one-hot encoder.
@@ D.1 Scenes to test evaluation model
D.2 Scenes to test conditional evaluation model
We collected the A-share data of China from 2024-8 to 2024-11 (89 days) by an open source tool AkShare. The collected day-wise features of stocks are “Profit Ratio”, “Average Cost”, “90 Cost Low”, “90 Cost High”, “90 Episode Medium”, “70 Cost Low”, “70 Cost High”, “70 Episode Medium”, “Closing”, “Opening”, “Highest”, “Lowest”, “Transaction Volume”, “Transaction Amount”, “Amplitude”. The strategy is to buy 1 hand, sell 1 hand, or hold by the agent following the open price of the market. The slippage and commission is not considered in this work. If the stock was limit up (limit down), then we can only hold even the model decision is to buy (sell). The agent’s input is the feature of the stock last day and output is the decision this day. The agent will be deployed and run for 1 time slot (10 trading days, about 2 weeks), and the evaluation condition is feature of the stock in the last time slot. At the end of deployment in a time slot, we will sell all the hold stocks. When the end of the slot is limit down at open, we assume that it will be sold by 90 percents of open price in future.
In order to simulate the direct effect from the agent itself to the sensitive market, we add a float model which input is the vectorized agents and output is float rate of open price ranged from -0.1% to +0.1%.
We use data of 200 stocks with minimal ID from all 3214 stocks due to our memory limit. The start of train day is from day 10 to day 41, the start of test day is from 41 to 61. The sampling number of agent is 200. The experiment was performed 5 times to reduce the influence of randomness. The proxy metric is the last 10 days’ RoI.
Appendix E Assumption assessment
Due to the limitation of computation resources and the difference of addressed problems, we do not use the deep learning models for observation data as base learners of our evaluation models, such as CEVAE (Louizos et al. [2017]), BNR (Li and Fu [2017]), BNN (Johansson et al. [2016]), SITE (Yao et al. [2018]), GANITE (Yoon et al. [2018]), DeepMatch (Kallus [2020]), Dragonnet (Shi et al. [2019]), TARNet/CFR (Shalit et al. [2017]), TabPFN (Hollmann et al. [2025]), deconfounder (Wang and and [2019]), time-series deconfounder (Bica et al. [2020]). Table LABEL:tab:assessregression and table LABEL:tab:assessclass are assumption assessment results of 7 kinds of evaluation models with swift base learners for continuous and discrete target respectively. From the tables, we can see that the IIDE assumption (IID check and ID check) and unbiased evaluation model assumption (Bias Check and GroupBias Check) was not violated obviously with a high probability (>0.8) except the cases that using MLP and SVR as base regression learners of HetEM.
Table 6: The ratio that the p-value of test is larger than or equal to 0.05 in the 30 repeated experiments for regression base learners. The name of each row is “EvaluationModelName”-“TestName”-“MetricName”. The name of each column is the scene name.
@@ Appendix E Assumption assessment
Het(Linear)-IID-RMSE 0.93 0.87 0.97 0.9 0.97
Het(Linear)-IID-$R^{2}$ 0.93 1.0 0.9 0.93 0.83
Het(Linear)-ID-RMSE 1.0 0.47 0.87 1.0 0.93
Het(Linear)-ID-$R^{2}$ 0.97 0.43 0.93 0.97 0.87
Het(Linear)-Bias-RMSE 0.9 0.43 0.77 0.8 0.77
Het(Linear)-Bias-$R^{2}$ 0.9 0.33 0.8 0.77 0.77
Het(Linear)-GroupBias-RMSE 0.92 0.85 0.87 0.87 0.85
Het(Linear)-GroupBias-$R^{2}$ 0.90 0.85 0.86 0.84 0.89
Het(MLP)-IID-RMSE 1.0 0.93 0.83 1.0 0.9
Het(MLP)-IID-$R^{2}$ 0.9 0.97 0.93 0.97 0.97
Het(MLP)-ID-RMSE 0.73 0.37 0.87 0.73 0.77
Het(MLP)-ID-$R^{2}$ 0.9 0.3 0.77 0.77 0.67
Het(MLP)-Bias-RMSE 0.5 0.26 0.57 0.57 0.5
Het(MLP)-Bias-$R^{2}$ 0.57 0.17 0.53 0.57 0.57
Het(MLP)-GroupBias-RMSE 0.68 0.40 0.86 0.70 0.72
Het(MLP)-GroupBias-$R^{2}$ 0.70 0.31 0.71 0.67 0.68
Het(SVR)-IID-RMSE 0.97 0.97 0.97 1.0 0.87
Het(SVR)-IID-$R^{2}$ 0.93 0.9 1.0 0.9 0.97
Het(SVR)-ID-RMSE 0.0 0.0 0.27 0.4 0.1
Het(SVR)-ID-$R^{2}$ 0.0 0.0 0.0 0.2 0.0
Het(SVR)-Bias-RMSE 0.0 0.0 0.1 0.17 0.03
Het(SVR)-Bias-$R^{2}$ 0.0 0.23 0.0 0.03 0.0
Het(SVR)-GroupBias-RMSE 0.0 0.01 0.19 0.34 0.11
Het(SVR)-GroupBias-$R^{2}$ 0.0 0.44 0.00 0.19 0.02
Het(RF)-IID-RMSE 0.97 0.93 0.97 0.97 0.93
Het(RF)-IID-$R^{2}$ 1.0 0.97 0.97 1.0 0.97
Het(RF)-ID-RMSE 0.93 0.87 0.9 0.97 0.83
Het(RF)-ID-$R^{2}$ 0.93 0.93 0.93 0.9 0.73
Het(RF)-Bias-RMSE 0.93 0.57 0.9 0.9 0.8
Het(RF)-Bias-$R^{2}$ 0.83 0.63 0.93 0.87 0.8
Het(RF)-GroupBias-RMSE 0.87 0.85 0.91 0.90 0.86
Het(RF)-GroupBias-$R^{2}$ 0.89 0.86 0.91 0.88 0.86
Het(LGBM)-IID-RMSE 0.87 0.93 0.97 1.0 0.93
Het(LGBM)-IID-$R^{2}$ 1.0 0.97 0.97 0.93 0.97
Het(LGBM)-ID-RMSE 0.93 0.9 0.9 0.97 0.83
Het(LGBM)-ID-$R^{2}$ 0.9 0.93 0.83 0.97 0.93
Het(LGBM)-Bias-RMSE 0.9 0.7 0.8 0.9 0.77
Het(LGBM)-Bias-$R^{2}$ 0.93 0.7 0.77 0.9 0.83
Het(LGBM)-GroupBias-RMSE 0.89 0.85 0.90 0.88 0.85
Het(LGBM)-GroupBias-$R^{2}$ 0.89 0.86 0.97 0.88 0.87
Het(XGBoost)-IID-RMSE 1.0 0.93 0.93 1.0 0.87
Het(XGBoost)-IID-$R^{2}$ 0.93 0.93 0.93 0.97 0.93
Het(XGBoost)-ID-RMSE 0.93 0.9 1.0 0.93 0.97
Het(XGBoost)-ID-$R^{2}$ 0.93 0.9 0.93 0.93 0.87
Het(XGBoost)-Bias-RMSE 0.97 0.53 0.77 0.87 0.87
Het(XGBoost)-Bias-$R^{2}$ 0.83 0.5 0.9 0.87 0.83
Het(XGBoost)-GroupBias-RMSE 0.92 0.84 0.83 0.86 0.88
Het(XGBoost)-GroupBias-$R^{2}$ 0.87 0.82 0.90 0.84 0.88
Het(CatBoost)-IID-RMSE 0.87 0.97 0.97 1.0 1.0
Het(CatBoost)-IID-$R^{2}$ 0.87 0.97 0.93 0.93 1.0
Het(CatBoost)-ID-RMSE 0.93 0.87 0.93 0.97 0.97
Het(CatBoost)-ID-$R^{2}$ 0.97 0.9 0.93 0.97 1.0
Het(CatBoost)-Bias-RMSE 0.93 0.6 0.97 0.8 0.83
Het(CatBoost)-Bias-$R^{2}$ 0.93 0.63 0.93 0.8 0.87
Het(CatBoost)-GroupBias-RMSE 0.90 0.90 0.92 0.87 0.87
Het(CatBoost)-GroupBias-$R^{2}$ 0.90 0.87 0.90 0.86 0.88
Table 7: The ratio that the p-value of test is larger than or equal to 0.05 in the 30 repeated experiments for classification base learners. The name of each row is “EvaluationModelName”-“TestName”-“MetricName”. The name of each column is the scene name.
@@
Het(Linear)-IID-ROCAUC 0.96 0.93 0.93 0.93 1.0 0.97
Het(Linear)-IID-ACC 0.93 0.97 0.9 0.9 0.97 0.97
Het(Linear)-ID-ROCAUC 1.0 0.97 0.87 0.9 1.0 0.97
Het(Linear)-ID-ACC 0.93 0.97 0.87 0.93 0.93 1.0
Het(Linear)-Bias-ROCAUC 0.77 0.83 0.83 0.9 0.9 0.9
Het(Linear)-Bias-ACC 0.83 0.87 0.87 0.8 0.83 0.93
Het(Linear)-GroupBias-ROCAUC 0.83 0.88 0.87 0.88 0.91 0.87
Het(Linear)-GroupBias-ACC 0.87 0.89 0.87 0.88 0.87 0.89
Het(MLP)-IID-ROCAUC 0.97 1.0 0.97 0.97 0.93 0.93
Het(MLP)-IID-ACC 0.93 0.93 0.97 0.9 0.97 0.97
Het(MLP)-ID-ROCAUC 0.8 0.87 0.77 0.87 0.87 0.7
Het(MLP)-ID-ACC 0.8 0.83 0.83 0.8 0.73 0.8
Het(MLP)-Bias-ROCAUC 0.43 0.57 0.53 0.53 0.63 0.47
Het(MLP)-Bias-ACC 0.67 0.6 0.57 0.57 0.43 0.6
Het(MLP)-GroupBias-ROCAUC 0.70 0.76 0.71 0.73 0.77 0.69
Het(MLP)-GroupBias-ACC 0.77 0.73 0.77 0.74 0.63 0.73
Het(SVM)-IID-ROCAUC 0.97 0.97 0.93 0.97 0.97 1.0
Het(SVM)-IID-ACC 0.97 0.97 0.97 0.97 1.0 1.0
Het(SVM)-ID-ROCAUC 0.73 0.5 0.9 0.6 0.83 0.8
Het(SVM)-ID-ACC 0.9 0.67 0.37 0.97 0.97 0.97
Het(SVM)-Bias-ROCAUC 0.4 0.27 0.1 0.37 0.53 0.47
Het(SVM)-Bias-ACC 0.97 0.37 0.43 0.93 0.8 0.93
Het(SVM)-GroupBias-ROCAUC 0.63 0.48 0.36 0.52 0.69 0.66
Het(SVM)-GroupBias-ACC 0.90 0.56 0.62 0.90 0.85 0.86
Het(RF)-IID-ROCAUC 0.97 0.97 0.9 0.97 0.97 0.97
Het(RF)-IID-ACC 0.93 0.97 0.87 0.93 0.9 0.9
Het(RF)-ID-ROCAUC 0.97 0.87 0.9 0.9 0.83 0.97
Het(RF)-ID-ACC 1.0 0.87 0.9 0.97 1.0 0.87
Het(RF)-Bias-ROCAUC 0.73 0.9 0.77 0.87 0.93 0.87
Het(RF)-Bias-ACC 0.9 0.93 0.8 0.97 0.8 0.7
Het(RF)-GroupBias-ROCAUC 0.85 0.87 0.84 0.9 0.91 0.86
Het(RF)-GroupBias-ACC 0.87 0.89 0.85 0.89 0.88 0.85
Het(LGBM)-IID-ROCAUC 1.0 0.97 0.97 0.97 0.87 1.0
Het(LGBM)-IID-ACC 0.93 0.97 0.93 0.9 1.0 0.9
Het(LGBM)-ID-ROCAUC 1.0 0.87 0.93 0.93 0.87 0.93
Het(LGBM)-ID-ACC 0.97 0.83 0.97 1.0 0.97 0.97
Het(LGBM)-Bias-ROCAUC 0.7 0.93 0.83 0.73 0.97 0.9
Het(LGBM)-Bias-ACC 0.83 0.9 0.77 0.83 0.8 0.8
Het(LGBM)-GroupBias-ROCAUC 0.81 0.88 0.85 0.89 0.91 0.86
Het(LGBM)-GroupBias-ACC 0.88 0.88 0.84 0.90 0.86 0.85
Het(XGBoost)-IID-ROCAUC 1.0 0.97 1.0 0.8 0.93 1.0
Het(XGBoost)-IID-ACC 0.93 1.0 0.97 1.0 0.93 0.87
Het(XGBoost)-ID-ROCAUC 0.9 0.83 0.9 0.9 0.87 1.0
Het(XGBoost)-ID-ACC 0.97 0.87 0.93 0.97 0.93 0.97
Het(XGBoost)-Bias-ROCAUC 0.83 0.97 0.83 0.9 0.83 0.83
Het(XGBoost)-Bias-ACC 0.87 0.93 0.9 0.83 0.8 0.87
Het(XGBoost)-GroupBias-ROCAUC 0.88 0.89 0.90 0.90 0.87 0.89
Het(XGBoost)-GroupBias-ACC 0.88 0.89 0.88 0.88 0.89 0.88
Het(CatBoost)-IID-ROCAUC 0.93 0.8 0.9 0.97 0.93 0.9
Het(CatBoost)-IID-ACC 1.0 0.9 1.0 0.87 1.0 0.97
Het(CatBoost)-ID-ROCAUC 0.97 0.93 0.93 0.93 0.87 0.97
Het(CatBoost)-ID-ACC 0.97 0.93 0.83 0.97 0.97 0.97
Het(CatBoost)-Bias-ROCAUC 0.83 0.87 0.8 0.8 0.9 0.87
Het(CatBoost)-Bias-ACC 0.87 0.9 0.93 0.87 0.8 0.8
Het(CatBoost)-GroupBias-ROCAUC 0.86 0.89 0.88 0.86 0.89 0.87
Het(CatBoost)-GroupBias-ACC 0.85 0.89 0.89 0.90 0.86 0.86
@@
We also visualization the Q-Q plot of the baseline evaluation model in figure 6 and the assumption assessment result of our evaluation model in table 8. From the visualization, the error distribution of baseline is close to normal distribution. Although the HetEM-Linear and HetEM-CatBoost does not pass our ID check and Bias check, the dramatically empirical error reduction of our evaluation model is still believed as valuable if upper bound existed. Of course it can be not used to calculate the upper bound of the two evaluation models by theorem 4 directly in this scene.
@@
Het(Linear)-IID 1.0
Het(Linear)-ID 0.0
Het(Linear)-Bias 0.0
Het(CatBoost)-IID 1.0
Het(CatBoost)-ID 0.0
Het(CatBoost)-Bias 0.0
Table 8: The ratio that p-value of test is larger than or equal to 0.05 in the 5 repeated experiments for A-share trade scene. The name of each row is “EvaluationModelName”-“TestName”-“MetricName”. The name of each column is the scene name.
Figure 6: Q-Q plot of evaluation error of Baseline (Last10DayRoI).
Appendix F Hyper-parameters of evaluation model